If the marginal utility of good 1 is 15 and its price is $3, and the marginal utility of good 2 is 12, and its price is $4, then the consumer___________

should increase his or her consumption of good 2.

should increase his or her consumption of good 1.

has achieved a consumer optimum

should decrease his or her consumption of good 1.
Answer #1

Answer is option B)

At eqm MU1/P1 = MU2/P2

now MU1/P1 = 15/3 = 5

MU2/P2 = 12/4 = 3

SO MU1/P1 > MU2/P2

SO MARGINAL UTILITY DERIVED FROM ADDITIONAL SPENDING ON GOOD 1 IS ALWAYS HIGHER THAN MU DERIVED FROM ADDITIONAL SPENDING ON GOOD 2

SO CONSUMER SHOULD INCREASE THE CONSUMPTION OF GOOD 1